Lily Tucker (single) owns and operates a bike shop as a sole proprietorship. In 2019, she sells the following long-term assetJust need the tax liability and how you got it!!
Answer #1
Calculation of tax liability
There are 2 components involved here :
1. Ordinary Income
2. Capital gains Income
Calculation of tax liability on Ordinary Income
Ordinary Income = $194,900
Tax Slab Rate Taxable Amount Tax liability
0 to 9700 10%                        9,700                970
9701 to 39475 12%                      29,775             3,573
39476 to 84200 22%                      44,725             9,840
84201 to 160725 24%                      78,325           18,798
160726 to 194900 32%                      34,175           10,936
Total tax liability on Ordinary Income :          44,117
Calculation of tax liability on Capital Gains
Capital Gains Income = $45,800
Capital Gains Tax rate for 2019 = 15%
Capital Gains tax = 45,800*15% = 6,870 $
HENCE, TOTAL TAX LIABILITY = 44,117+6,870 = 50,987$

Note/Assumptions :

1. Since it is not mentioned in the questions as to which year's tax rates/slabs are to be used, we have used the 2019 federal income tax rates for the calculation of tax liability. We have also mentioned the rates used below :

2019 federal income tax brackets

Tax rate Single
10% $0 to $9,700
12% $9,701 to $39,475
22% $39,476 to $84,200
24% $84,201 to $160,725
32% $160,726 to $204,100
35% $204,101 to $510,300
37% $510,301 or more

2. In absence of any information, we have assumed this gain to be long-term capital gain i.e., asset is sold after 1 year and hence applied the tax rate of 15%. Had these been short term capital gains, normal tax rates would have been applied.